The distribution of scores for the 1,000 final exams in a statistics course has a population mean of 74 and a population standard deviation of 15. A random sample of 36 exam papers is selected. What is the probability that the sample mean is higher than 77? (a) 0.1100 (b) 0.2151 (c) 1131 (d)1151
Answer #1

P (* >77) = P(- 74 77-74 = P( Z > 3/2.5) = P(Z > 1.2 ) = 1 – P(Z < 1.2) = 1 - 0 .88493 = 0.1151

Where Z is a standard normal variable

So the correct answer is option d.

PLEASE LET ME KNOW IF YOU HAVE ANY DOUBTS. THANKS!!